# Glauber predictions for oxygen and neon collisions at the LHC

### v2

@@ VII Particle production
The centrality dependence of charged-particle production in heavy-ion collisions was found to scale with the number of participants PHOBOS:2010eyu . In particular at LHC energies Loizides:2011ys , inclusion of a small contribution from the number of collisions improves the description to account for the increase of multiplicity in more central collisions from mini-jets, as ${\rm d}N/{\rm d}\eta=\alpha N_{\rm part}+\beta N_{\rm coll}$. Typically, ${\rm d}N/{\rm d}\eta$ is normalized per participant pair, $N_{\rm part}/2$. When this formalism is applied across different collision energies and system sizes, using sub-nucleon degrees of freedom —specifically, three or more “constituent” quarks— makes it unnecessary to rely on the number of binary nucleon-nucleon collisions since one can achieve approximate scaling with the generalized number of constituent participants (see Loizides:2016djv and references therein).
@@
Here, we take a different approach, and introduce the number of multiple hard interactions as implemented in HIJING Wang:1991hta and applied recently in Loizides:2017sqq ; Jonas:2021xju . For a given nucleon–nucleon collision the number of hard scatterings at $b_{\rm NN}$ is distributed according to a Poisson distribution with the average given by $\langle N^{\rm mpi}_{\rm NN}\rangle=\sigma_{\rm NN}^{\rm hard}\,T_{\rm NN}(b_{\rm NN})$, where $\sigma_{\rm NN}^{\rm hard}$ is the energy-dependent pQCD cross-section for $2\to 2$ parton scatterings ($\approx 190$ mb at 5.36 TeV). The total number of multiple interactions $N_{\rm mpi}$ is then the sum of the multiple interactions over all nucleon–nucleon collision in the event, and on average equal to $N_{\rm coll}$ times the average $N_{\rm mpi}$ in pp collisions at the same energy. Since the $N_{\rm mpi}$ and event multiplicity are strongly correlated Loizides:2021ima , ordering events according to $N_{\rm mpi}$ can be used at a proxy for event categorization usually done to group events into centrality intervals. In this case, the relation on the average between $N_{\rm mpi}$ and $N_{\rm coll}$ does no longer need to hold. The left panel of Fig. 12 shows the resulting $\left\langle N_{\rm mpi}\right\rangle$ distributions normalized by $\left\langle N_{\rm part}\right\rangle$ in centrality-binned intervals versus $N_{\rm part}$ for OO, XeXe and PbPb collisions at 5.36 TeV. After a steep rise at very low $N_{\rm part}$, the $N_{\rm mpi}$ in the different systems rise approximately linearly with $N_{\rm part}$, until they eventually increase faster than linear with $N_{\rm part}$, most evidently seen for OO and XeXe collisions. In the right panel, we show $R\equiv\left\langle N_{\rm mpi}\right\rangle/(\left\langle N_{\rm coll}\right\rangle\,N_{\rm mpi}^{\rm pp})$ versus $N_{\rm part}$, which exhibits the effect of the geometry and event selection bias on the nuclear modification factor, even in absence of any nuclear effect Loizides:2017sqq . Indeed, the ratio significantly deviates from one in peripheral collisions, where the ratio is smaller up to about 50%, and while it is enhanced by up to 10% in the most central OO collisions. In reality, effects from fragmentation, momentum conservation, play a role, which are expected to result in a $p_{\rm T}$-dependent selection bias Park:2025mbt .
Per-participant charged particle pseudo-rapidity density $2/\left\langle N_{\rm part}\right\rangle\,{\rm d}N/{\rm d}\eta$ versus $\left\langle N_{\rm part}\right\rangle$ at mid-rapidity in XeXe collisions at 5.44 TeV and in PbPb collisions at 5.36 TeV obtained by ALICE and CMS ALICE:2018cpu ; CMS:2019gzk ; CMS:2024ykx ; ALICE:2025cjn are shown in Fig. 13. In case of XeXe collisions the ALICE and CMS date are not consistent for peripheral collisions. Data for pp collisions normalized to non-single diffractive and inelastic yields are also shown (scaled to 5.36 TeV by less than one percent as given by respective power laws as a function of $\sqrt{s}$ALICE:2022kol .
@@
Figure 13: Per-participant charged particle pseudo-rapidity density $2/\left\langle N_{\rm part}\right\rangle\,{\rm d}N/{\rm d}\eta$ versus $\left\langle N_{\rm part}\right\rangle$ at mid-rapidity in XeXe collisions at 5.44 TeV (left) and in PbPb collisions at 5.36 TeV (right) obtained by ALICE and CMS ALICE:2018cpu ; CMS:2019gzk ; CMS:2024ykx ; ALICE:2025cjn , as well as two component fits to each dataset (see text). Data ALICE:2022kol for pp collisions normalized to non-single diffractive and inelastic yields are also shown.
@@
Figure 14: Per-participant charged particle pseudo-rapidity density $2/\left\langle N_{\rm part}\right\rangle\,{\rm d}N/{\rm d}\eta$ versus $\left\langle N_{\rm part}\right\rangle$ at mid-rapidity in XeXe collisions at 5.44 TeV and in PbPb collisions at 5.36 TeV obtained by ALICE and CMS ALICE:2018cpu ; CMS:2019gzk ; CMS:2024ykx ; ALICE:2025cjn , as well as data for pp collisions normalized to non-single diffractive and inelastic yields ALICE:2022kol . The figures zooms into the low $N_{\rm part}$ region to show highlight predictions for OO and NeNe collisions at 5.36  TeV obtained by average of the fits obtained in Fig. 13 together with the respective two-component model fits (see text).
@@
While a disagreeing trend is visible in the PbPb case, the dataset are consistent considering the large uncertainties in peripheral collisions. Since the data are essentially at the same collision energy as that for the planned OO collisions, they can be used to predict the ${\rm d}N/{\rm d}\eta$ for OO collisions, by fitting a two-component model $2N(f+(1-f)\,{\left\langle N_{\rm mpi}\right\rangle}/{\left\langle N_{\rm part}\right\rangle})$. The fits are found to describe the data reasonably well, however with different values for $N$ between XeXe and PbPb collisions, in particular for ALICE. We hence assigned as predicted result the central value between the ratios and half of the spread as uncertainty to essentially cover all cases. The same procedure was also applied to give the expected mid-rapidity ${\rm d}N/{\rm d}\eta$ for NeNe collisions. The expectations for OO and NeNe collisions at 5.36 TeV are shown in Fig. 14 together with the existing data. A table with all relevant Glauber related values and projected multiplicities is given in Tab. 5. The Glauber values are given without uncertainties. However, it should be noted that the associated uncertainties usually estimated from data require careful evaluation, and are expected to be non-negligible.
@@ VIII Summary
An updated version (v3.3) of the TGlauberMC model is presented, which includes several different probability distributions for the nucleon–nucleon interaction probability (Fig. 1), and a huge variety of nuclear density profiles (Tab. 1 and 2), in particular for oxygen and neon (Fig. 2 and 3). The model is used to predict properties of OO, NeNe, as well as pPb and PbPb collisions at 5.36 TeV, and pO collisions at 9.62 TeV. The expected geometrical cross sections are averaged over all provided profiles (Fig. 4), and considered to include a contribution from the transverse nucleon–nucleon overlap (Fig. 5 and  6), and presented in Tab. 4. For central collisions, the distributions of $N_{\rm coll}$ and $N_{\rm part}$ exhibit a spread of 5–10% and about 2%, respectively (Fig. 7), and up to 20% for the participant eccentricity (Fig. 8). The ratio of $\varepsilon_{2}$ in NeNe to OO is found to reach up to $1.15\pm 0.05$ (Fig. 9). Rising trends for $\varepsilon_{3}$ and $\varepsilon_{4}$ with decreasing centrality (Fig. 10), while first decreasing then rising trends with decreasing centrality of the respective ratios between NeNe and OO collisions (Fig. 11) are found. Using $N_{\rm mpi}$ instead of $N_{\rm coll}$ (Fig. 12) allows to predict general trends of the OO data (Fig. 13) and use existing data at 5.36 TeV to predict the multiplicity in OO and NeNe collisions (Fig. 14). The source code for TGlauberMC (v3.3) is available at HepForge glaucode , and the author welcomes suggestions how to improve it.
@@ Acknowledgements.
The author would like to thank G. Nijs and W. van der Schee for fruitful discussions related to nuclei configurations provided by Trajectum, G. Giacalone for clarification related to the 3pF fit of the NNLOsat oxygen profile, V. Soma for providing the exact data points of the NNLOsat calculation Soma:2019bso , B. Kardan for contributing changes in TGlauberMC v3.3 useful for Hades/CBM collision energies, as well as F. Jonas and N. Strangmann for taking the time to read and comment on a draft of this article. The author acknowledges financial support by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under contract number DE-SC0005131.
@@
The main changes related to the previous version (3.2) is the interface to any nucleus configuration provided by Trajectum Nijs:2021clz via the new TrNucGen trnucgen library (see App. B), the support to read nucleus configurations from text files, and a larger list pre-defined nucleus configurations. Further, the possibility to use between overlap functions using TRENTO, HIJING and PYTHIA parameterizations, in addition to the $\Gamma$-distribution that already was present since v3.0. As default, the code continues to use the hard-sphere approximation ($\omega=0$, see Sec. III). Following the discussion in Sec. VII, the total number of multiple interactions are computed from the nucleon–nucleon overlap function. The $\sigma_{\rm NN}$ parameterization for RHIC and LHC energies is provided, as well as dedicated parameterizations Bystricky:1987yq for neutron–neutron (proton–proton) and neutron–proton interactions relevant at Hades/CBM energies.
The user’s guide has been incrementally updated from the first release of the code in 2008 Alver:2008aq , the improved version in 2014 Loizides:2014vua , and the latest in 2017 Loizides:2017ack . As before, the source code, which relies on the ROOT Brun:1997pa framework (version 6.x), can be obtained at the TGlauberMC page on HepForge glaucode . For version 3.3, it also provides a doxygen documentation, which provides many details. In the following, we only describe the improved functionality relative to the last update Loizides:2017ack .
@@
where the position ($x$, $y$, $z$) and charge $c$ are listed per line for a nucleus configuration. The format of the file is also explained in the example provided with TrNucGen trnucgen ).
@@
The following set of functions controls additional behavior of the TGlauberMC class: SetOmega($w$) to provide the nucleon–nucleon profile to use. The input value can be $0<w<2$ for the $\Gamma$-distribution, $7$ for HIJING, $8$ for PYTHIA and $9<w<11$, which would transate into $0<\delta<2$ for TRENTO. SetSigmaHard to set the hard scattering cross section (in case you do not want to rely on the precomputed value.
@@
Additional macros that are provided are
@@ Appendix B TrNucGen
The Trajectum Nucleus Generator (TrNucGen) software library trnucgen provides all nuclei configurations available in Trajectum v2.1 Nijs:2021clz . It was written in particular to use all the configurations for Oxygen-16 and Neon-20 nuclei Giacalone:2024luz explored in this paper. The library is defines a convenient interface to the relevant code extracted from Trajectum, avoiding having to compile the latter with all the required dependencies. The full list of available oxygen and neon profiles from Trajectum is given in Tab. 6. The list of all nuclei configurations can be found in the doxygen documentation of the code. The installation instructions as well as the source code including a standalone example are publicly available trnucgen . The library can be also very useful for any heavy-ion model or simulation that would like to explore nuclei configurations beyond those based on common parameterizations (i.e. 3pF).
Table 6: Full list of oxygen-16 and neon-20 NLEFT and PGCM calculations taken from Trajectum Giacalone:2024luz .
@@
TR_OXYGEN_V1 Variational Monte Carlo version Lim:2018huo (same as “O” in TGlauberMC v2.7 and higher)
@@
TR_OXYGEN_V3 3pF from Summerfield:2021oex (same as “Opar” in TGlauberMC v2.7 and higher)
TR_OXYGEN_V4 3pF fit to NNLOSat Soma:2019bso (same as “Opar2” in TGlauberMC v3.3)

### v3

@@ VII Particle production
The centrality dependence of charged-particle production in heavy-ion collisions was found to scale with the number of participants Alver and others (2011). In particular at LHC energies Loizides (2011), inclusion of a small contribution from the number of collisions improves the description to account for the increase of multiplicity in more central collisions from mini-jets, as ${\rm d}N/{\rm d}\eta=\alpha N_{\rm part}+\beta N_{\rm coll}$. Typically, ${\rm d}N/{\rm d}\eta$ is normalized per participant pair, $N_{\rm part}/2$.
When this formalism is applied across different collision energies and system sizes, employing sub-nucleon degrees of freedom by decomposing each nucleon into three or more effective quark-level interaction centers—removes the need to reference the number of binary nucleon–nucleon collisions, as approximate scaling arises from the generalized number of participant quark centers (see Loizides (2016b) and references therein).
@@
Here, I take a different approach, and introduce the number of multiple hard interactions as implemented in HIJING Wang and Gyulassy (1991) and applied recently in Loizides and Morsch (2017); Jonas and Loizides (2021). For a given nucleon–nucleon collision the number of hard scatterings at $b_{\rm NN}$ is distributed according to a Poisson distribution with the average given by $\langle N^{\rm mpi}_{\rm NN}\rangle=\sigma_{\rm NN}^{\rm hard}\,T_{\rm NN}(b_{\rm NN})$, where $\sigma_{\rm NN}^{\rm hard}$ is the energy-dependent pQCD cross-section for $2\to 2$ parton scatterings ($\approx 190$ mb at 5.36 TeV). The total number of multiple interactions $N_{\rm mpi}$ is then the sum of the multiple interactions over all nucleon–nucleon collision in the event, and on average equal to $N_{\rm coll}$ times the average $N_{\rm mpi}$ in pp collisions at the same energy. Since the $N_{\rm mpi}$ and event multiplicity are strongly correlated Loizides and Morsch (2021), ordering events according to $N_{\rm mpi}$ can be used at a proxy for event categorization usually done to group events into centrality intervals. In this case, the relation on the average between $N_{\rm mpi}$ and $N_{\rm coll}$ does no longer need to hold. The left panel of Fig. 13 shows the resulting $\left\langle N_{\rm mpi}\right\rangle$ distributions normalized by $\left\langle N_{\rm part}\right\rangle$ in centrality-binned intervals versus $N_{\rm part}$ for OO, XeXe and PbPb collisions at 5.36 TeV. After a steep rise at very low $N_{\rm part}$, the $N_{\rm mpi}$ in the different systems rise approximately linearly with $N_{\rm part}$, until they eventually increase faster than linear with $N_{\rm part}$, most evidently seen for OO and XeXe collisions. In the right panel, $R\equiv\left\langle N_{\rm mpi}\right\rangle/(\left\langle N_{\rm coll}\right\rangle\,N_{\rm mpi}^{\rm pp})$ versus $N_{\rm part}$ is shown, which exhibits the effect of the geometry and event selection bias on the nuclear modification factor, even in absence of any nuclear effect Loizides and Morsch (2017). Indeed, the ratio significantly deviates from one in peripheral collisions, where the ratio is smaller up to about 50%, and while it is enhanced by up to 10% in the most central OO collisions. In reality, effects from fragmentation, momentum conservation, play a role, which are expected to result in a $p_{\rm T}$-dependent selection bias Park et al. (2025).
@@
Figure 14: Per-participant charged particle pseudo-rapidity density $2/\left\langle N_{\rm part}\right\rangle\,{\rm d}N/{\rm d}\eta$ versus $\left\langle N_{\rm part}\right\rangle$ at mid-rapidity in XeXe collisions at 5.44 TeV (left) and in PbPb collisions at 5.36 TeV (right) obtained by ALICE and CMS Acharya and others (2019b); Sirunyan and others (2019); Hayrapetyan and others (2025d); Acharya and others (2025), as well as two component fits to each dataset (see text). Data Acharya and others (2023) for pp collisions normalized to non-single diffractive and inelastic yields are also shown.
@@
Per-participant charged particle pseudo-rapidity density $2/\left\langle N_{\rm part}\right\rangle\,{\rm d}N/{\rm d}\eta$ versus $\left\langle N_{\rm part}\right\rangle$ at mid-rapidity in XeXe collisions at 5.44 TeV and in PbPb collisions at 5.36 TeV obtained by ALICE and CMS Acharya and others (2019b); Sirunyan and others (2019); Hayrapetyan and others (2025d); Acharya and others (2025) are shown in Fig. 14. In case of XeXe collisions the ALICE and CMS date are not consistent for peripheral collisions. Data for pp collisions normalized to non-single diffractive and inelastic yields are also shown (scaled to 5.36 TeV by less than one percent as given by respective power laws as a function of $\sqrt{s}$Acharya and others (2023).
Figure 15: Per-participant charged particle pseudo-rapidity density $2/\left\langle N_{\rm part}\right\rangle\,{\rm d}N/{\rm d}\eta$ versus $\left\langle N_{\rm part}\right\rangle$ at mid-rapidity in XeXe collisions at 5.44 TeV and in PbPb collisions at 5.36 TeV obtained by ALICE and CMS Acharya and others (2019b); Sirunyan and others (2019); Hayrapetyan and others (2025d); Acharya and others (2025), as well as data for pp collisions normalized to non-single diffractive and inelastic yields Acharya and others (2023). The figures zooms into the low $N_{\rm part}$ region to show highlight predictions for OO and NeNe collisions at 5.36  TeV obtained by average of the fits obtained in Fig. 14 together with the respective two-component model fits (see text).
@@
Despite the apparent opposite trend in the PbPb case the dataset are consistent considering the large uncertainties in peripheral collisions. Since the data are essentially at the same collision energy as that for the planned OO collisions, they can be used to predict the ${\rm d}N/{\rm d}\eta$ for OO collisions, by fitting a two-component model $2N(f+(1-f)\,{\left\langle N_{\rm mpi}\right\rangle}/{\left\langle N_{\rm part}\right\rangle})$. The fits are found to describe the data reasonably well, however with different values for $N$ between XeXe and PbPb collisions, in particular for ALICE. Hence, for the predicted result, the central value between the ratios is used with half of the spread as uncertainty assigned to essentially cover all cases. The same procedure was also applied to give the expected mid-rapidity ${\rm d}N/{\rm d}\eta$ for NeNe collisions. The expectations for OO and NeNe collisions at 5.36 TeV are shown in Fig. 15 together with the existing data.
@@
A table with all relevant Glauber related values and projected multiplicities is given in Tab. 5. The Glauber values are given without uncertainties. However, it should be noted that the associated uncertainties usually estimated from data require careful evaluation, and are expected to be non-negligible.
@@ VIII Summary
An updated version (v3.3) of the TGlauberMC model is presented, which includes several different probability distributions for the nucleon–nucleon interaction probability (Fig. 1), and a huge variety of nuclear density profiles (Tab. 1 and 2), in particular for oxygen and neon (Fig. 2 and 3). The model is used to predict properties of OO, NeNe, as well as pPb and PbPb collisions at 5.36 TeV, and pO collisions at 9.62 TeV. The expected geometrical cross sections are averaged over all provided profiles (Fig. 4), and considered to include a contribution from the transverse nucleon–nucleon overlap (Fig. 5 and  6), and presented in Tab. 4. For central collisions, the distributions of $N_{\rm coll}$ and $N_{\rm part}$ exhibit a spread of 5–10% and about 2%, respectively (Fig. 7), and up to 20% for the participant eccentricity (Fig. 8). The ratio of $\varepsilon_{2}$ in NeNe to OO is found to reach up to $1.15\pm 0.05$ (Fig. 9), rather independently whether Gaussian smearing (Fig. 10) using generalized TRENTO conditions is used or not. Rising trends for $\varepsilon_{3}$ and $\varepsilon_{4}$ with decreasing centrality (Fig. 11), while first decreasing then rising trends with decreasing centrality of the respective ratios between NeNe and OO collisions (Fig. 12) are found. Using $N_{\rm mpi}$ instead of $N_{\rm coll}$ (Fig. 13) allows to predict general trends of the OO data (Fig. 14) and use existing data at 5.36 TeV to predict the multiplicity in OO and NeNe collisions (Fig. 15). The source code for TGlauberMC (v3.3) is available at HepForge 92, and the author welcomes suggestions how to improve it.
During the peer-review of the manuscript results from OO collisions at $\sqrt{s_{\rm NN}}=200$ GeV, taken in June, 2021 at RHICAboona and others (2025), and from OO and NeNe collisions at $5.36$ TeV at LHC I. J. Abualrob et al. (2025); G. Aad et al. (2025); A. Hayrapetyan et al. (2025b); 69; A. Hayrapetyan et al. (2025a); 34 were submitted. The availability of the OO data at both RHIC and LHC, opens an important opportunity to explore both particle production and collective flow across a wide energy range — and TGlauberMC (v3.3) should prove a valuable tool for interpreting those measurements.
@@ Acknowledgements.
The author would like to thank G. Nijs and W. van der Schee for fruitful discussions related to nuclei configurations provided by Trajectum, G. Giacalone for clarification related to the 3pF fit of the NNLOsat oxygen profile, V. Soma for providing the exact data points of the NNLOsat calculation Somà et al. (2020), B. Kardan for contributing changes in TGlauberMC v3.3 useful for Hades/CBM collision energies, as well as F. Jonas and N. Strangmann for taking the time to read and comment on a draft of this article. The author acknowledges financial support by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under contract number DE-SC0005131.
@@
The main changes related to the previous version (3.2) is the interface to any nucleus configuration provided by Trajectum Nijs and van der Schee (2022b) via the new TrNucGen 93 library (see App. B), the support to read nucleus configurations from text files, and a larger list pre-defined nucleus configurations. Further, the possibility to use between overlap functions using TRENTO, HIJING and PYTHIA parameterizations, in addition to the $\Gamma$-distribution that already was present since v3.0. As default, the code continues to use the hard-sphere approximation ($\omega=0$, see Sec. III). Following the discussion in Sec. VII, the total number of multiple interactions are computed from the nucleon–nucleon overlap function. The $\sigma_{\rm NN}$ parameterization for RHIC and LHC energies is provided, as well as dedicated parameterizations Bystricky et al. (1987) for neutron–neutron (proton–proton) and neutron–proton interactions relevant at Hades/CBM energies.
The user’s guide has been incrementally updated from the first release of the code in 2008 Alver et al. (2008), the improved version in 2014 Loizides et al. (2015), and the latest in 2017 Loizides et al. (2018). As before, the source code, which relies on the ROOT Brun and Rademakers (1997) framework (version 6.x), can be obtained at the TGlauberMC page on HepForge 92. For version 3.3, it also provides a “doxygen” documentation, which provides many details. In the following, I only describe the improved functionality relative to the last update Loizides et al. (2018).
@@
where the position ($x$, $y$, $z$) and charge $c$ are listed per line for a nucleus configuration. The format of the file is also explained in the example provided with TrNucGen 93).
@@
The following set of member functions controls additional behavior of the TGlauberMC class: SetOmega($w$) to provide the nucleon–nucleon profile to use. The input value can be $0<w<2$ for the $\Gamma$-distribution, $7$ for HIJING, $8$ for PYTHIA and $9<w<11$, which would translate into $0<\delta<2$ for TRENTO. SetSigmaHard to set the hard scattering cross section (in case you do not want to rely on the precomputed value. SetShadowing to control whether or not shadowing corrections are computed.
@@
Additional global functions that are provided are
@@
These also exists as corresponding static functions of the TGlauberMC class. Additional static functions are ReadNtuple to read the produced Ntuple, GetCentralityDist to produce a centrality distribution based on impact parameter and AddCentralityBranch to add the centrality estimator to the produced tree.
The runAndSmearNtuple() function has been improved to include generalized overlap function defined as $\left(\frac{(T_{A}^{p}+T_{B}^{p})}{2}\right)^{\frac{1}{p}}$ Moreland et al. (2015), which for $p=1$ reduces to the arithmetic mean and for $p=0$ to the geometric mean (also known as TRENTO initial conditions). The function produces a sample of $n$ events with a generalized reduced-thickness (or $p$-parameter) formulation of the overlap. The parameter $p$ controls the functional form of the participant thickness; unless stated otherwise the default value $p=1$ is used, while $p=0$ corresponds to the TRENTO prescription. The nuclei involved in the collision are specified by the strings sysA and sysB, which are interpreted through the TGlauNucleus::Lookup tables. The inelastic nucleon–nucleon cross section is set by $\sigma_{NN}$ (in units of mb), for which a default value of $68\,\mathrm{mb}$ is used. A minimum nucleon–nucleon separation $d_{\min}$ is enforced during nucleon placement, with $d_{\min}=0.4\,\mathrm{fm}$ in our baseline configuration. The model allows for a tunable nucleon profile characterized by the parameter $\omega$. A value $\omega=0$ corresponds to a hard-sphere profile, $0\leq\omega\leq 2$ yields a Gamma-distributed profile, and $\omega=7$ and $8$ reproduce the effective profiles used in HIJING and PYTHIA, respectively. Larger values ($\omega=9$$11$) correspond to the generalized TRENTO implementations. Additional profile-shape parameters include $w$, for which the TRENTO-motivated value $w=0.5$ is used, as well as $k$, the shape parameter of the Gamma distribution, with default value $k=1.4$. Events are sampled within an impact-parameter range $b_{\min}\leq b\leq b_{\max}$, where $b_{\min}=0$ and $b_{\max}=20\,\mathrm{fm}$ are used unless otherwise stated. The resulting events are written to an output ntuple specified by fname; if no file name is supplied, a default name is automatically generated.
@@ Appendix B TrNucGen
The Trajectum Nucleus Generator (TrNucGen) software library 93 provides all nuclei configurations available in Trajectum v2.1 Nijs and van der Schee (2022b). It was written in particular to use all the configurations for Oxygen-16 and Neon-20 nuclei Giacalone and others (2025) explored in this paper. The library is defines a convenient interface to the relevant code extracted from Trajectum, avoiding having to compile the latter with all the required dependencies. The full list of available oxygen and neon profiles from Trajectum is given in Tab. 6. The list of all nuclei configurations can be found in the doxygen documentation of the code. The installation instructions as well as the source code including a standalone example are publicly available 93. The library can be also very useful for any heavy-ion model or simulation that would like to explore nuclei configurations beyond those based on common parameterizations (i.e. 3pF).
Table 6: Full list of oxygen-16 and neon-20 NLEFT and PGCM calculations taken from Trajectum Giacalone and others (2025).
@@
TR_OXYGEN_V1 Variational Monte Carlo version Lim et al. (2019) (same as “O” in TGlauberMC v2.7 and higher)
@@
TR_OXYGEN_V3 3pF from Summerfield et al. (2021) (same as “Opar” in TGlauberMC v2.7 and higher)
TR_OXYGEN_V4 3pF fit to NNLOSat Somà et al. (2020) (same as “Opar2” in TGlauberMC v3.3)